OpenStudy (anonymous):

What is the slope of the line through the points (-2, 7) and (3, -3)? m = _?_

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) =(-3-7)/(3-(-2)) =-10/5 =-2

OpenStudy (anonymous):

M=(rise)/(run) = 10/-5 = -2